# Robust Spatial Confounding Adjustment via Basis Voting

### v1

@@ 4.2 Drop-One Estimator
As seen in the previous section is that the projection based method estimator, although is consistent for large samples has a finite-sample bias $\frac{\bm{h}_{n,j}^{\top}\bm{U}_{n}}{\bm{h}_{n,j}^{\top}\bm{X}_{n}}$. Even when $j\in O_{X}$, and the term vanishes asymptotically, it still contributes to the asymptotic variance of $\hat{\beta}_{j,n}^{PROJ}$. We present an improved candidate estimator by including basis functions from the set $\mathcal{H}_{U}$ in Figure 1. These variables support $U$ but not $X$ and are thus not part of the confounding mechanism. They are often termed as “precision variables” in the causal literature, as including them in the regression helps explain more of the outcome variance, thereby increasing precision of the estimator. For $j\in\{1,\cdots,d\}$, define the matrix $\bm{H}_{-j|1:d}=\begin{bmatrix}\bm{h}_{n,1}&\cdots&\bm{h}_{n,j-1}&\bm{h}_{n,j+1}&\cdots&\bm{h}_{n,d}\end{bmatrix}$ i.e. the first $d$ basis evaluations dropping the $j$-th one. We can decompose $U(s)=\sum_{l=1}^{\infty}\alpha_{u}^{(l)}h_{l}(s)=\alpha_{u}^{(j)}h_{j}(s)+\sum_{l\neq j}^{d}\alpha_{u}^{(l)}h_{l}(s)+(1-\lambda_{d})U_{>d}(s)$ for $1\leq j\leq d$, where $\lambda_{d}=1$ if $U$ has support only on the first $d$ basis functions and 0 otherwise and hence $U_{>d}(s)=\sum_{l=d+1}^{\infty}\alpha_{u}^{(l)}h_{l}(s)$. Hence we can write $\bm{U}_{n}=\alpha_{u}^{(j)}\bm{h}_{j,n}+\bm{H}_{-j|1:d}\bm{\alpha}_{-j,u}+(1-\lambda_{d})\bm{U}_{n,>d}$ where $\bm{\alpha}_{-j,u}=\left(\alpha_{u}^{(1)},\cdots,\alpha_{u}^{(j-1)},\alpha_{u}^{(j+1)},\cdots,\alpha_{u}^{(d)}\right)^{\top}$. If $j\in O_{X}$, this simplifies to $\bm{U}_{n}=\bm{H}_{-j|1:d}\bm{\alpha}_{-j,u}+(1-\lambda_{d})\bm{U}_{n,>d}$. Thus, the DGP for the observed values projected onto the first $d$ basis functions can be written as:
@@
Thus if $j\in O_{X}$, the second term above is 0. Also, as the columns of $\bm{P}_{d}$ are asymptotically uncorrelated with $\bm{U}_{n,>d}$, $\bm{P}_{d}\bm{U}_{n,>d}/n$ vanishes asymptotically. Hence, for $j\in O_{X}$, we can write the regression as
$\mathbb{E}\left[\frac{1}{n}\bm{Y}_{n,d}\right]\approx\frac{1}{n}\tilde{\bm{X}}_{-j}\bm{\gamma}_{-j,u},$
@@
This candidate estimator is similar in spirit to the residual-on-residual regression proposed by Thaden & Kneib (2018), although we only use it as a candidate estimator fed into our ensemble mode estimator of Section 3, thereby allowing us to relax the assumption of rough exposure as considered there. The drop-one candidate estimator incorporates more information about the data into the regression of $\bm{Y}_{n,d}$ on $\tilde{\bm{X}}_{-j}$ as compared to the projection method, where we were choosing one basis function at a time, and hence the former should be statistically more efficient. The following theorem shows the performance of the drop-one estimator in large samples:
@@
Under the DGP as in Equation 1 and the estimator as in Equation 7, assuming a sampling density $f_{S}>0$ on the compact domain $\mathcal{D}$ for the i.i.d. sampled coordinates with $\epsilon(S_{i})|S_{i}$’s being iid with mean 0 and variance $\sigma^{2}_{\epsilon}$, defining $r_{j,d}=\alpha_{u}^{(j)}/\alpha_{x}^{(j)}$ such that $j\in S_{x,d}$, we have:
@@
    $\hat{\beta}_{-j,n}\xrightarrow[n\to\infty]{P}\beta+r_{j,d}$.
@@
    For any $j\in O_{X}$, we have $\sqrt{n}\left(\hat{\beta}_{-j,n}-\beta\right)\xrightarrow[n\to\infty]{d}N\left(0,\frac{(1-\lambda_{d})\sigma^{2}_{U,j,>d}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}\right)$ where $\sigma^{2}_{U,j,>d}\coloneqq Var(h_{j}U_{>d})=\|h_{j}U_{>d}\|^{2}_{f_{S}}$.
@@
    If $\lambda_{d}=1$, then for $j\in O_{X}$ we have $\mathbb{E}\left[\hat{\beta}_{-j,n}\right]=\beta$.
The first two results are similar to those in Theorem 2, establishing consistency and asymptotic normality of the candidate drop-one estimators. Additionally, if $U$ admits a finite basis expansion with respect to $\mathcal{H}$ which is implictly assumed by many existing spatial deconfounding approaches, e.g., the spline-based adjustment of Keller & Szpiro (2020), the third result shows that additionally $\hat{\beta}_{-j,n}$ is exactly finite-sample unbiased for any $j\in O_{X}$. This unbiasedness property is an improvement over the projection-based estimator which always has a finite-sample bias.
4.3 Comparison of Drop-one and Projection Method
We now look at the comparison of asymptotic variances of the estimators based on valid candidate basis functions which support $X$ but not $U$, i.e., $h_{j}$ for $j\in O_{X}$. We have seen from Theorems 2 and 3 that for any $j\in O_{X}$, the candidate’s estimates are consistent for $\beta$ and asymptotically normal with mean $\beta$. Also, we can see the asymptotic variance of the scaled drop-one estimator is $avar\left(\hat{\beta}_{-j,n}\right)=\dfrac{(1-\lambda_{d})\sigma^{2}_{U,j,>d}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}$ and that for the projection estimator is $avar\left(\hat{\beta}^{PROJ}_{j,n}\right)=\frac{\sigma^{2}_{U,j}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}$. Once again, under an assumed finite basis representation of $U$, the term $(1-\lambda_{d})\sigma^{2}_{U,j,>d}$ vanishes and because $\sigma^{2}_{U,j}>0$ as discussed at the end of section 4.1, we immediately have that:
@@ 4.3 Comparison of Drop-one and Projection Method
i.e., our proposed drop-one candidate is more efficient than the projection candidate.
The following lemma shows that the result holds even if $U$ has infinite expansion for many choices of basis functions.
Lemma 2.
Let $\mathcal{H}=\{h_{j}\}_{j\geq 1}$ be orthonormal in $\mathcal{L}^{4}(\mathcal{D},f_{S})$, with $U=\sum_{j\geq 1}\alpha_{u}^{(j)}h_{j}$, $U_{\leq d}=\sum_{j\leq d}\alpha_{u}^{(j)}h_{j}$, $U_{>d}=\sum_{j>d}\alpha_{u}^{(j)}h_{j}$ for $d\in\mathbb{N}$ and define $S_{\mathrm{in}}\coloneqq\sum_{j\leq d}{\alpha_{u}^{(j)}}^{2}$. Assume:
  1. (i)
    $\max_{1\leq j\leq d}\|h_{j}\|_{\infty}\leq M<\infty$.
  2. (ii)
  3. (iii)
Then for every $j=1,\dots,d$ with $j\in O_{X}$ i.e. $\alpha_{u}^{(j)}=0$,
$\|h_{j}U\|^{2}_{f_{S}}>\|h_{j}U_{>d}\|^{2}_{f_{S}}.$
Using this lemma, it can be directly inferred from Theorems 2 and 3 that $\sigma^{2}_{U,j}>\sigma^{2}_{U,j,>d}$ for $j\in O_{X}$. Thus even for $U$ having possibly infinite expansion i.e. for $\lambda_{d}=0$, if $\mathcal{H}$ satisfies the assumptions of Lemma 2, then:
The first two assumptions of Lemma 2 (where $\|h_{j}\|_{\infty}\coloneqq\sup_{s}|h_{j}(s)|$) are not restrictive and are satisfied by most of the standard families of basis functions used for analysis in a compact spatial domain, such as Fourier basis and polynomial basis classes. The third assumption is not specific to any basis class and is satisfied for a choice of $d$ that is sufficiently large. The fact that $c^{\ast}>0$ can be derived from Assumption (ii) of the lemma, more details about which can be found in the Appendix. This shows that the drop-one candidate estimator is more efficient than the standard projection-based estimator under a broad range of scenarios, motivating its use in the ensemble mode estimator of Section 3.
@@ 5.1 Simulation Data Generation
To investigate the performance of the proposed estimator under controlled conditions, we have conducted a comprehensive simulation study. The primary goal of the simulation was to evaluate whether basis voting could recover the causal effect $\beta$ in the presence of unmeasured spatial confounding, when the structure of confounding aligns with our identification assumptions, which may or may not match with that of the existing methods.
@@
The non-zero values of the exposure coefficients vector $\left\{\alpha_{x}^{(j)}\right\}$ and confounder coefficients vector $\left\{\alpha_{u}^{(j)}\right\}$ were sampled independently from $Unif(3,6)$ distribution along with random sign $\{-1,1\}$ assigned with equal probability. The coefficients were kept fixed over all the data replicates as we are in a deterministic function scenario. The magnitude of the confounding bias in naive OLS was calibrated based on the inner product of the coefficient vectors, ensuring a challenging identification task.
@@ 5.2 Correct Basis Specification
We simulate two distinct scenarios for the exposure $X(s)$ and the unmeasured confounder $U(s)$ to evaluate the performance of our proposed method relative to existing approaches. Specifically, we consider cases where (a) the exposure is a rougher function of space than the confounder, and (b) the exposure is smoother. Many existing methods assume the exposure is spatially rougher than the confounder, and we aim to assess the robustness of our method when this assumption is violated.
Figure 4: Performance of different methods for estimating $\beta=2.5$ under two scenarios: (a) when the exposure $X$ is rougher than the unmeasured confounder $U$ and (b) when $X$ is smoother than $U$, using correct basis specification. Our mode-based estimators (named Proj-One Mode and Drop-One Mode) recover the effect sizes in both cases; competitors bias when $X$ is smoother.
Figure 4 compares our proposed estimator to several existing methods, using the oracle estimator as a benchmark (which leverages knowledge of the true shared basis functions for optimal efficiency). Competing methods include Ordinary Least Squares (OLS), Generalized Least Squares (GLS) with a pre-specified exponential covariance kernel using the BRISC R-package (Saha & Datta 2018), the spatial basis adjustment approach using TPRS basis set in a similar essence to Keller & Szpiro (2020) by adding a spline term in the regression (to adjust for the unmeasured confounder $U$) additional to the exposure of interest, the gSEM method of Thaden & Kneib (2018), the Spatial+ approach of Dupont et al. (2022), and the spectral domain adjustment of Guan et al. (2023). The results reveal important differences in method performance. When the exposure is rougher than the unmeasured confounder with respect to the Fourier basis (Panel (a)), nearly all methods, including ours, accurately recover the true value of $\beta=2.5$. Only exception is the unadjusted OLS estimator which of course is expected to be biased under confounding. When the exposure is smoother than the confounder (Panel (b)), all existing methods yield substantially biased estimates, while our mode-based estimators — both project-on-one and drop-one approaches – remain unbiased and accurately estimate $\beta$. This experiment highlights the robustness of our method to violation of the rough exposure assumption.
@@ 5.3 Robustness to Basis Misspecification
The experiments in the previous section were in the idealized scenario in which the basis functions used for estimation match those that generated the data. Note that using the correct basis is not a requirement for our method, which only requires that the plurality holds under whichever basis set is used in the analysis to represent the exposure and the confounder. Here, we assess the robustness of our method to basis misspecification.
@@
As we see, in both scenarios, the coefficient ratios have the biggest clustering around the $y=0$ line, indicating that the modal value remains close to zero. This supports the validity of Assumption 2 for this class of basis functions, regardless of whether the exposure is smoother or rougher. Similar findings were obtained for the eigen basis functions.
Figure 6 presents the estimated effect sizes produced by our method when using the first $d$ basis functions (with $d$ varying from 150 to 850, for a sample size of 900 observations). Results are presented for both TPRS and eigen basis sets, and under both relative smoothness scenarios for $X$ and $U$.
@@
These results show that even under basis misspecification, stable estimation of the effect size is attained once a sufficient number of basis functions are included in the model for the plurality rule to kick in. In terms of bias, the projection method performs slightly worse as compared to the drop-one methods in finite samples. Also the projection method exhibits a greater standard error, consistent with the theoretical findings in Section 4.3. We can see that as we go on increasing the number of basis functions in the model, the standard error of the estimates also increases, which is meaningful as we are using more and more variables in the model, which should yield in higher standard error of the estimates, which shows that there is a bias-variance tradeoff here. A higher number of basis functions in the model can definitely help reduce the bias of the estimate, but at the cost of increasing the standard error. Overall, the drop-one candidates based mode estimator is much more robust to the number of basis functions used. The experiments in this section show that our mode estimator works well even with a misspecified basis, both when the exposure is rougher or smoother, and that using the drop-one candidate estimators generally leads to better performance than the projection-based candidate estimators.

### v2

@@ 4.2 Drop-One Estimator
Despite being consistent, the projection-based estimator has finite-sample bias $\frac{\bm{h}_{n,j}^{\top}\bm{U}_{n}}{\bm{h}_{n,j}^{\top}\bm{X}_{n}}$. Even when $j\in O_{X}$ and the term vanishes asymptotically, it still contributes to the asymptotic variance of $\hat{\beta}_{j,n}^{PROJ}$. We can improve the efficiency of the candidate estimator by including basis functions from the set $\mathcal{H}_{U}$ (see right panel of Figure 1). These variables support $U$ but not $X$ and are thus not part of the confounding mechanism. Instead, they are precision variables: predictors of the outcome but not exposure, the inclusion of which increases the precision of the regression by explaining some of the variation in the outcome that is not due to or confounded with the exposure.
We first provide the intuition behind the approach. For simplicity, suppose $U$ admits a finite basis expansion up to $d$ bases. For any $j\in\{1,2,\ldots,d\}$ which supports $X$ but not $U$, i.e., $j\in O_{X,d}$, we drop the $j^{th}$ basis function and regress the outcome $Y$ on $X$ and $h_{1},h_{2},..h_{j-1},h_{j+1},\ldots,h_{d}$. This adjusts for all confounders, i.e., all $h_{k}$ which support both $X$ and $U$, and also includes as precision variables all $h_{k^{\prime}}$ that support only $U$ but not $X$. Leaving out $h_{j}$ ensures that there is residual variation left in $X$ that is not accounted for the included set of basis functions, adjusting for the confounders ensures that the coefficient corresponding to $X$ is consistent for $\beta$ leveraging this residual variation in $X$ and the linearity of the regression, and including precision variables reduces the variance of this candidate compared to the projection-based estimator.
We now describe the process formally without assuming a true finite basis expansion of $U$. Let $\bm{H}_{1:d}=\begin{bmatrix}\bm{h}_{n,1}&\cdots&\bm{h}_{n,d}\end{bmatrix}$ be the evaluation matrix for the first $d$ basis functions and denote the corresponding projection on its column space by $\bm{P}_{d}=\bm{H}_{1:d}\left(\bm{H}_{1:d}^{\top}\bm{H}_{1:d}\right)^{-1}\bm{H}_{1:d}^{\top}$. Define the projected exposures and outcomes on the first $d$ basis as $\bm{X}_{n,d}=\bm{P}_{d}\bm{X}_{n}$ and $\bm{Y}_{n,d}=\bm{P}_{d}\bm{Y}_{n}$. For $j\in\{1,\cdots,d\}$, define the matrix $\bm{H}_{-j|1:d}=\begin{bmatrix}\bm{h}_{n,1}&\cdots&\bm{h}_{n,j-1}&\bm{h}_{n,j+1}&\cdots&\bm{h}_{n,d}\end{bmatrix}$, i.e., the first $d$ basis functions, dropping the $j$-th one. We can decompose $U(s)=\sum_{l=1}^{\infty}\alpha_{u}^{(l)}h_{l}(s)=\alpha_{u}^{(j)}h_{j}(s)+\sum_{l\in 1:d,l\neq j}\alpha_{u}^{(l)}h_{l}(s)+(1-\lambda_{d})U_{>d}(s)$ for $1\leq j\leq d$, where $\lambda_{d}=1$ if $U$ has support only on the first $d$ basis functions and 0 otherwise and hence $U_{>d}(s)=\sum_{l=d+1}^{\infty}\alpha_{u}^{(l)}h_{l}(s)$. Hence we can write $\bm{U}_{n}=\alpha_{u}^{(j)}\bm{h}_{n,j}+\bm{H}_{-j|1:d}\bm{\alpha}_{-j,u}+(1-\lambda_{d})\bm{U}_{n,>d}$ where $\bm{\alpha}_{-j,u}=\left(\alpha_{u}^{(1)},\cdots,\alpha_{u}^{(j-1)},\alpha_{u}^{(j+1)},\cdots,\alpha_{u}^{(d)}\right)^{\top}$. Thus, the DGP for the observed values projected onto the first $d$ basis functions can be written as:
@@
If $j\in O_{X,d}$, the second term above is 0. Also, as the columns of $\bm{P}_{d}$ are asymptotically uncorrelated with $\bm{U}_{n,>d}$, $\bm{P}_{d}\bm{U}_{n,>d}/n$ vanishes asymptotically. Hence, for $j\in O_{X,d}$, we can write the regression as
$\mathbb{E}\left[\bm{Y}_{n,d}\right]\approx\tilde{\bm{X}}_{-j}\bm{\gamma}_{-j,u},$
@@
The following theorem formalizes our claim that by adjusting for $\bm{H}_{-j|1:d}$, the drop-one candidate estimator both adjusts for confounders and accounts for variation in the outcome due to the precision variables and is therefore a more efficient consistent estimator than the projection-based estimator.
@@
Consider the DGP as in Equation 1 and the estimator as in Equation 7, assuming a sampling density $f_{S}>0$ on the compact domain $\mathcal{D}$ for the i.i.d. sampled coordinates with $\epsilon(S_{i})|S_{i}$’s being iid with mean 0 and variance $\sigma^{2}_{\epsilon}$. Let $d$ be such that $O_{X,d}$ is non-empty. Defining $r_{j}=\alpha_{u}^{(j)}/\alpha_{x}^{(j)}$ such that $j\in S_{X,d}$, we have:
@@
    $\hat{\beta}_{-j,n}\xrightarrow[n\to\infty]{P}\beta+r_{j}$.
@@
    For any $j\in O_{X,d}$, we have $\sqrt{n}\left(\hat{\beta}_{-j,n}-\beta\right)\xrightarrow[n\to\infty]{d}N\left(0,\frac{(1-\lambda_{d})\sigma^{2}_{U,j,>d}+\sigma^{2}_{\epsilon}}{{\alpha_{x}^{(j)}}^{2}}\right)$ where $\sigma^{2}_{U,j,>d}\coloneqq Var(h_{j}U_{>d})=\|h_{j}U_{>d}\|^{2}_{f_{S}}$.
@@
    If $\lambda_{d}=1$, then for $j\in O_{X,d}$ we have $\mathbb{E}\left[\hat{\beta}_{-j,n}\right]=\beta$.
The first two results are similar to those in Theorem 2, establishing consistency and asymptotic normality of the candidate drop-one estimators. Additionally, if $U$ admits a finite basis expansion with respect to $\mathcal{H}$, which is implicitly assumed by many existing spatial deconfounding approaches, e.g., the spline-based adjustment of keller2020selecting, the third result shows that additionally $\hat{\beta}_{-j,n}$ is exactly finite-sample unbiased for any $j\in O_{X,d}$. This unbiasedness property is an improvement over the projection-based estimator, which has a finite-sample bias.
4.3 Comparison of Drop-one and Projection-based Candidates
We now look at the comparison of asymptotic variances of the estimators based on valid candidate basis functions which support $X$ but not $U$, i.e., $h_{j}$ for $j\in O_{X}$. For a finite sample comparison, see the simulations in Section 5.3. We have seen from Theorems 2 and 3 that for any $j\in O_{X,d}$ respectively, both the candidate’s estimates are consistent for $\beta$ and asymptotically normal with mean $\beta$. The following theorem formally states that the asymptotic variance of the drop-one estimator is smaller than that of the projection estimator:
Theorem 4.
Consider the DGP as in Equation 1 and the estimator as in Equation 7, assuming a sampling density $f_{S}>0$ on the compact domain $\mathcal{D}$ for the i.i.d. sampled coordinates with $\epsilon(S_{i})|S_{i}$’s being iid with mean 0 and variance $\sigma^{2}_{\epsilon}$. Let the basis functions $h_{1},h_{2},\ldots,$ be such that each $h_{j}$ has a zero set $\{s\in\mathcal{D}:h_{j}(s)=0\}$ of $f_{S}$-measure zero. Define the candidate estimators $\hat{\beta}^{PROJ}_{j,n}$ and $\hat{\beta}_{-j,n}$ as in Equations 5 and 7 respectively based on basis functions $h_{1},\ldots,h_{d}$. Then for large enough $d$ and $j\in O_{X,d}$,
@@ 4.3 Comparison of Drop-one and Projection-based Candidates
where $avar(\cdot)$ denotes the asymptotic variance.
The assumption of Theorem 4, that the zero set of each of the basis functions are of measure zero, is satisfied by many of the standard families of basis functions used for analysis in a compact spatial domain with a suitable choice of the sampling density $f_{S}>0$, such as Fourier basis and polynomial basis classes. The results proves that the drop-one candidate estimator is more efficient than the standard projection-based candidate estimator under a broad range of scenarios. Therefore, we recommend using the drop-one candidates in the basis-voting estimation procedure of Section 3.
@@ 5.1 Simulation Data Generation
To investigate the performance of the proposed estimator under controlled conditions, we have conducted a comprehensive simulation study. The primary goal of the simulation was to evaluate whether basis voting could recover the causal effect $\beta$ in the presence of unmeasured spatial confounding, when the structure of confounding aligns with our plurality assumption.
@@
The non-zero values of the exposure coefficients vector $\left\{\alpha_{x}^{(j)}\right\}$ and confounder coefficients vector $\left\{\alpha_{u}^{(j)}\right\}$ were sampled independently from $Unif(3,6)$ distribution along with random sign $\{-1,1\}$ assigned with equal probability. The coefficients were kept fixed over all the data replicates. The magnitude of the confounding bias in naive OLS was calibrated based on the inner product of the coefficient vectors, ensuring a challenging identification task.
@@ 5.2 Correct Basis Specification
We simulate two distinct scenarios: (a) the exposure is a rougher function of space than the confounder, and (b) the exposure is smoother. Many existing methods assume the exposure is spatially rougher than the confounder, and we aim to assess the robustness of all methods when this assumption is violated.
Figure 4: Performance of different methods for estimating $\beta=2.5$ under two scenarios: (a) when the exposure $X$ is rougher than the unmeasured confounder $U$ and (b) when $X$ is smoother than $U$, using the correct basis specification. Our mode-based estimators (named Proj-One Mode and Drop-One Mode) recover the effect sizes in both cases; competitors bias when $X$ is smoother.
Figure 4 compares our proposed estimator to several existing methods, using the oracle estimator as a benchmark (which leverages knowledge of the true $U$ and the shared basis functions for optimal efficiency). Competing methods include Ordinary Least Squares (OLS), Generalized Least Squares (GLS) with a pre-specified exponential covariance kernel using the BRISC R-package (saha2018brisc), the Spatial basis adjustment approach using TPRS basis set in a similar essence to keller2020selecting by adding a spline term in the regression (to adjust for the unmeasured confounder $U$) additional to the exposure of interest, the gSEM method of thaden2018structural, the Spatial+ approach of dupont2022spatial+, and the Standard Matérn model-based spectral domain adjustment of guan2023spectral. The results reveal important differences in method performance. When the exposure is rougher than the unmeasured confounder with respect to the Fourier basis (Panel (a)), nearly all methods, including ours, accurately recover the true value of $\beta=2.5$. The only exception is the unadjusted OLS estimator, which of course is expected to be biased under confounding. When the exposure is smoother than the confounder (Panel (b)), all existing methods yield substantially biased estimates, while our mode-based estimators — both project-on-one and drop-one approaches – are almost unbiased and accurately estimate $\beta$. This experiment highlights the robustness of our method to violation of the rough exposure assumption.
@@ 5.3 Robustness to Basis Misspecification
The experiments in the previous section were in the idealized scenario in which the basis functions used for estimation match those that generated the data. Using the correct basis is not a requirement for our method, which only requires that the plurality holds under whichever basis set is used in the analysis to represent the exposure and the confounder. Here, we assess the robustness of our method to basis misspecification.
@@
As we see, in both scenarios, the coefficient ratios have the biggest clustering around the $y=0$ line, indicating that the modal value remains close to zero. This supports the validity of Assumption 2 for this class of basis functions which was not used to generate the data, regardless of whether the exposure is smoother or rougher. Similar findings were obtained for the eigen basis functions.
Figure 6 presents the estimated effect sizes produced by our method when using the first $d$ basis functions (with $d$ varying from 150 to 850, for a sample size of 900 observations), as the true value of $d_{0}$ from Assumption 2 is unknown, especially under basis misspecification. Results are presented for both TPRS and eigen basis sets, and under both relative smoothness scenarios for $X$ and $U$. By varying $d$ from small values up to nearly $n$, we observe that the estimator stabilizes as the number of candidate basis estimates increases and the plurality rule becomes more clearly satisfied. Evaluating the estimator across a range of $d$ therefore provides a practical safeguard against the instability that can arise when $d$ is chosen to be too small for the plurality condition to hold.
@@
These results show that even under basis misspecification, stable estimation of the effect size is attained once a sufficient number of basis functions are included in the model for the plurality rule to kick in. In terms of bias, the projection method performs worse as compared to the drop-one methods in finite samples. Also, the projection method exhibits a greater standard error. Both of these trends are consistent with the theoretical findings in Section 4.3. We can see that as the number of basis functions in the model increases, the standard error of the estimates also increases, illustrating a bias-variance tradeoff in the number of basis functions included: more basis functions will generally reduce the bias of the estimate, but at the cost of increasing the standard error due to more parameters being estimated. Overall, the drop-one candidates based mode estimator is much more robust to the number of basis functions used. The experiments in this section show that our mode estimator works well even with a misspecified basis, both when the exposure is rougher or smoother, and that using the drop-one candidate estimators generally leads to better performance than the projection-based candidate estimators.